the 2019 bar exams only 2103 of the 7685 bar passed. The passing a weighted average of 74, which is higher than the mean weighted average of all the bar examinees. Find this mean given that the weighted average is normally distributed with a standard deviation of 16.62. Round off the final answer to the nearest whole number
1
Expert's answer
2020-10-19T17:43:27-0400

Let "X=" the weighted average: "X\sim N(\mu, \sigma^2)."

Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0,1)"


"P(X>74)=1-P(X\leq74)="

"=1-P(Z\leq\dfrac{74-\mu}{16.62})=\dfrac{2103}{7685}"

"P(Z\leq\dfrac{74-\mu}{16.62})=\dfrac{5582}{7685}\approx0.72635"

"\dfrac{74-\mu}{16.62}\approx0.601811"

"\mu\approx64"